OpenStudy (geometry_hater):

How do i do this quadratic equation? 6x squared + 7x - 10 = -5 gotta use the quadratic formula but what it that?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

Go down to where it says quadratic formula. you use that to solve for x

OpenStudy (anonymous):

but you have to have it in the form of ax^2 + bx + c= 0 you have 6x^2 + 7x -10 = -5 6x^2 + 7x -5 = 0

OpenStudy (anonymous):

You can enter enter 6x^2 + 7x -5 into your calculator and look for where y=0